OpenStudy (1018):

Help with matrix problem : find the determinant given that A isa 3x3 matrix for which det(A)=-3 (a) det(3A^T) (b) det(2A^-1)

OpenStudy (anonymous):

Okay, det(A) = -3, what's det(A^T) ?

OpenStudy (1018):

No, that's the problem. Those are the only given.

OpenStudy (1018):

I'm confused on what's supposed to do. lol

OpenStudy (anonymous):

Yes, but you should know these two properties by heart, ok? Read carefully... \[\large \det(A^T) = \det(A)\][Transposes have the same determinant] \[\large \det(A^{-1}) = \frac1{\det(A)}\]

OpenStudy (1018):

So, does it mean that that the answer is just for (a) is 3(-3)?

OpenStudy (1018):

Or am I getting it wrong?

OpenStudy (anonymous):

No... it isn't that simple, I'm afraid... but it IS still rather simple... Here's another property: \[\Large \det(cA) = c^n\det(A)\] Where n is the dimension of the matrix A.

OpenStudy (1018):

So how do I actually do it? Can you show me an example. Help me with (a).

OpenStudy (anonymous):

Certainly :) \[\Large \det(3A^T) = 3^3 \det(A^T)\] right?

OpenStudy (1018):

Wait, so the size of the matrix is 3x3, right? So how do I get 3 as the exponent?

OpenStudy (anonymous):

BECAUSE it's a 3x3 matrix :) If it was 4x4, the exponent would be 4 :D

OpenStudy (1018):

I mean, how to solve for it, like you get it. Haha

OpenStudy (1018):

I have a question, what if it's a 2x3? Haha

OpenStudy (anonymous):

You're being silly... non-square matrixes don't even have determinants :D

OpenStudy (1018):

Oooh. I don't know that. Hahaha

OpenStudy (1018):

Ok, back to the problem.

OpenStudy (anonymous):

\[\Large \det(3A^T) = 3^3 \det(A^T)\] You can work from here, you know the determinant of A^T, and you know 3^3

OpenStudy (1018):

Ok thanks! Sorry man for late replies, I'm running errands. thanks!